# Robust Policy Switching for Antifragile Reinforcement Learning for UAV Deconfliction in Adversarial Environments

### v1

@@ Appendix B Sampling Based Solution for Robust RL
$\hat{g_{s}}\left(\Theta_{t}\right)\leftarrow\nabla_{\Theta}\hat{\mathcal{L}}_{s}\left(\Theta_{t}\right).$ (B.1)
Here, ^$\hat{\mathcal{L}}_{s}\left(\Theta_{t}\right)$ represents the stochastic estimate of the loss function. Hence, the update equation can be represented as:
$\Theta_{t+1}\leftarrow\Theta_{t}-\eta\left(\hat{g_{s}}\left(\Theta_{t}\right)\right).$ (B.2)
@@
$\begin{gathered}\zeta_{t}\sim\mathcal{N}\left(0,\mathfrak{e}\right),\\ \Theta_{t+1}\leftarrow\Theta_{t}-\eta\left(\hat{g}_{s}\left(\Theta_{t}\right)+\zeta_{t}\right).\end{gathered}$ (B.3)
@@
$\begin{gathered}\bm{C}_{t}\leftarrow\rho\bm{C}_{t-1}+\left(1-\rho\right)\left(\hat{\bm{g}_{s}}\left(\Theta_{t}\right)-\bm{\mu}_{t}\right)\left(\hat{\bm{g}_{s}}\left(\Theta_{t}\right)-\bm{\mu}_{t-1}\right),\\ \zeta_{t}\sim\mathcal{N}\left(\bm{\mu}_{t},\bm{C}_{t}\right),\\ \Theta_{t+1}\leftarrow\Theta_{t}-\eta\left(\hat{\bm{g}}_{s}\left(\Theta_{t}\right)+\psi\zeta_{t}\right).\end{gathered}$ (B.4)
In equation (B.4), we observe that the noise covariance pre-conditioner, scales the noise proportionally along the dimensions with larger gradients. This approach helps to escape saddle points and navigate the policy space more effectively while dealing with broader loss minima, ultimately leading to faster convergence and improved solutions. Therefore, incorporating adaptive noise into the loss function allows optimization of both the critic loss and the agent and adversarial policy loss functions. If $\alpha_{i}$ as the adversarial contribution, then the general action $a_{t}$ is expressed as follows:
$a_{t}=\alpha_{i}\nu_{\omega}\left(\Phi_{t}\right)+\left(1-\alpha_{i}\right)\pi_{\theta}\left(\Phi_{t}\right).$ (B.5)
@@
$y_{\textup{targ}}=r_{t}+\gamma\left(1-\mathfrak{d}\right)Q_{\phi_{\textup{targ}}}\left(\Phi^{\prime},\left(1-\alpha_{i}\right)\mu_{\theta}+\alpha_{i}\nu_{\omega}\right).$ (B.6)
@@
$L\left(\phi\right)=\frac{1}{N}\sum_{B=\left\{\left(\Phi_{t},a_{t},r_{t},{\Phi}^{\prime},\mathfrak{d}\right)\right\}}\left(y\left(r,\Phi^{\prime},d\right)-Q_{\phi}\left(\phi,a\right)\right)^{2}.$ (B.7)
@@
$\begin{gathered}\nabla_{\theta}\widehat{J\left(\theta,\omega_{t}\right)}=\frac{1-\alpha_{i}}{N}\sum_{\Phi\in\mathcal{D}}\nabla_{\theta}\mu_{\theta}\left(\Phi\right)\nabla_{a}Q_{\phi}\left(\Phi,a\right),\\ \nabla_{\omega}\widehat{J\left(\theta_{t},\omega\right)}=\frac{\alpha_{i}}{N}\sum_{\Phi\in\mathcal{D}}\nabla_{\omega}\nu_{\omega}\left(\Phi\right)\nabla_{a}Q_{\phi}\left(\Phi,a\right).\end{gathered}$ (B.8)
To minimize the loss functions of critics and policy given by equations (B.6) and (B.7) respectively, we incorporate SGLD noise into the update equations as specified in (B.4). This approach is taken to update the joint critic parameters $\phi$ and the agent and adversary policy parameters $\left\{\theta,\omega\right\}$. If we consider $\bm{g}=\left\{\nabla_{\theta}J,\nabla_{\omega}J\right\}$, $\bm{\mu}_{t}=\left\{\mu_{\theta},\mu_{\omega}\right\}$, $\bm{C}_{t}=\left\{C_{\theta},C_{\omega}\right\}$ according to the dynamics of SGLD given in (B.4), the update equation for the agent and the adversary parameter to obtain the mixed Nash equilibrium for the max-min objective in (10) for the $k^{\textup{th}}$ sample from the buffer at time $t$, is given as:
$\theta^{k+1}_{t}\leftarrow\theta^{k}_{t}+\eta\left(\nabla_{\theta}\widehat{J\left(\theta,\omega_{t}\right)}+\psi\xi_{t}\right),$ (B.9)
$\omega^{k+1}_{t}\leftarrow\omega^{k}_{t}-\eta\left(\nabla_{\omega}\widehat{J\left(\theta,\omega_{t}\right)}+\psi\xi_{t}\right).$ (B.10)
@@
$\bar{\omega}_{t}=\left(1-\beta\right)\bar{\omega}_{t}+\beta\omega^{\left(k+1\right)}_{t};\;\bar{\theta}_{t}=\left(1-\beta\right)\bar{\theta}_{t}+\beta\theta^{\left(k+1\right)}_{t}.$ (B.11)
@@
$\omega_{t+1}=\left(1-\beta\right)\omega_{t}+\beta\bar{\omega}_{t};\;\theta_{t+1}=\left(1-\beta\right)\theta_{t}+\beta\bar{\theta}_{t}.$ (B.12)
@@
1:Observe state $\Phi$, select the actions according to $\rho_{\textup{agent}}=\mu_{\theta}+\zeta$, $\rho_{\textup{adv}}=\nu_{\omega}+{\zeta}^{\prime}$ where $\zeta$,${\zeta}^{\prime}\sim\mathcal{N}\left(0,\sigma I\right)$.
@@
9:         Sample a random minibatch of $N$ transitions $B=\left\{\left(\Phi,\bar{\rho},r,{\Phi}^{\prime},d\right)\right\}$ from $\mathcal{D}$.
10:         Compute the targets as per (B.6).
11:         Compute the loss function to update the joint critic $\nabla_{\phi}L\left(\phi\right)$ as per (B.7).
12:         Compute the agent and adversary policy gradient according to (B.8).
13:         Compute the agent and adversary parameters as per (B.8).
14:         Update the parameters of the agent and adversary network $\left\{\theta,\omega\right\}$ according to (B.9) and (B.10).
15:         Obtain the value of $\left\{\bar{\theta},\bar{\omega}\right\}$ according to (B.11).
16:         Update the parameters of the target network used to compute the target according to (B.9).
$\begin{gathered}\phi_{\textup{targ}}\leftarrow\tau\phi_{\textup{targ}}+\left(1-\tau\right)\phi\\ \theta_{\textup{targ}}\leftarrow\tau\theta_{\textup{targ}}+\left(1-\tau\right)\theta^{k+1}_{t}\\ \omega_{\textup{targ}}\leftarrow\tau\omega_{\textup{targ}}+\left(1-\tau\right)\omega^{k+1}_{t}\end{gathered}$ (B.13)
@@
18:     Update $\left\{\omega_{t+1},\theta_{t+1}\right\}$ as per (B.12).
@@
[…]
@@
$\begin{gathered}\alpha_{i}\leftarrow\alpha_{i}+0.1,\\ \mathcal{P}^{\prime}\leftarrow\mathcal{P}\cup\left\{\theta,\omega\right\}_{\alpha_{i}},\\ \mathcal{Q}^{\prime}\leftarrow\mathcal{Q}\cup\left\{\phi\right\}_{\alpha_{i}},\\ \mathcal{P}\leftarrow\mathcal{P}^{\prime},\\ \mathcal{Q}^{\prime}\leftarrow\mathcal{Q}.\end{gathered}$ (B.14)
@@
Appendix C Algorithm for Generating Adversarial States
Algorithm 2 Generate adversarial state $\Phi^{\textup{adv}}$ for obtaining the distributional shift from vanilla and robust policies
@@ Appendix C Algorithm for Generating Adversarial States
4:     Compute the step size ϝ$\digamma=c\mathbin{/}k+c$.
@@
Appendix D Algorithm for Discounted Thompson Sampling
@@
Appendix E Benchmarking Algorithms
Benchmarking is performed by evaluating the standard policy gradient algorithm, such as the DDPG, along with its robust variants. An approach involves an agent who intends to execute an action $\mathbf{a}$, but an alternative adversarial action $\bar{\mathbf{a}}$ is implemented with probability $\alpha$. Another variant, conceptually similar to the approach described in Section E.II, employs soft probabilistic robust policy iteration (PR-PI) as proposed in [tessler2019action], rather than solving the mixed Nash equilibrium objective in (10). The final benchmarking strategy involves training the agent to optimize the rewards without directly modifying the policy during training. Instead, an adversarial buffer is utilized, allowing the agent to learn the optimal value function in the presence of adversarial states. A brief explanation of these algorithms is provided in the subsequent subsections.
E.I Deep Deterministic Policy Gradient (DDPG)
The schematic shown in Figure 3, the action robust RL is replaced by DDPG considering $\alpha=0$. Hence, the critic target $y_{\textup{targ}}$ in (B.6) is modified as follows,
$y_{\textup{targ}}=r_{t}+\gamma\left(1-d\right)Q_{\phi_{\textup{targ}}}\left(\Phi^{\prime},\mu_{\theta}\right).$ (E.1)
Similarly the policy loss with $\alpha=0$ shown in (B.8) to compute the policy parameter $\theta$, will be as,
$\nabla_{\theta}\widehat{J\left(\theta\right)}=\frac{1}{N}\sum_{\Phi\in\mathcal{D}}\nabla_{\theta}\mu_{\theta}\left(\Phi\right)\nabla_{a}Q_{\phi}\left(\Phi,a\right).$ (E.2)
@@ E.I Deep Deterministic Policy Gradient (DDPG)
E.II Probabilistic Robust Markov Decision Process
@@
$\pi^{\textup{mix}}_{P,\alpha}\left(\left.\begin{matrix}\mathbf{a}\end{matrix}\right|\Phi\right)\equiv\left(1-\alpha\right)\pi\left(\left.\begin{matrix}\mathbf{a}\end{matrix}\right|\Phi\right)+\alpha\bar{\pi}\left(\left.\begin{matrix}\mathbf{a}\end{matrix}\right|\Phi\right).$ (E.3)
@@
E.III Noisy Robust Markov Decision Process
@@
$\pi^{\textup{mix}}_{N,\alpha}\left(\left.\begin{matrix}\mathbf{a}\end{matrix}\right|\Phi\right)=\mathbb{E}_{\mathbf{b}\sim\pi\left(.|\Phi\right),\bar{\mathbf{b}}\sim\bar{\pi}\left(.|\Phi\right)}\left[\vec{\mathbf{1}}_{\mathbf{a}=\left(1-\alpha\right)\mathbf{b}+\alpha\bar{\mathbf{b}}}\right].$ (E.4)
The NR-MDP is analogous to the Markov decision process (MDP) used to develop the ensembles, except for the addition of SGLD noise to the loss function, as described in equation (B.4) to obtain the mixed Nash equilibrium solution for the objective shown in (10). The solution to the min-max objective considered here is according to the soft PR-PI method as described in Algorithm 2 in [tessler2019action].
E.IV Learning with Adversarial Attacks

### v3

@@ Appendix D Sampling Based Solution for Robust RL
$\hat{g_{s}}\left(\Theta_{t}\right)\leftarrow\nabla_{\Theta}\hat{\mathcal{L}}_{s}\left(\Theta_{t}\right).$ (D.1)
Here, $\hat{\mathcal{L}}_{s}\left(\Theta_{t}\right)$ represents the stochastic estimate of the loss function. Hence, the update equation can be represented as:
$\Theta_{t+1}\leftarrow\Theta_{t}-\eta\left(\hat{g_{s}}\left(\Theta_{t}\right)\right).$ (D.2)
@@
$\begin{gathered}\zeta_{t}\sim\mathcal{N}\left(0,\mathfrak{e}\right),\\ \Theta_{t+1}\leftarrow\Theta_{t}-\eta\left(\hat{g}_{s}\left(\Theta_{t}\right)+\zeta_{t}\right).\end{gathered}$ (D.3)
@@
$\begin{gathered}\bm{C}_{t}\leftarrow\rho\bm{C}_{t-1}+\left(1-\rho\right)\left(\hat{\bm{g}_{s}}\left(\Theta_{t}\right)-\bm{\mu}_{t}\right)\left(\hat{\bm{g}_{s}}\left(\Theta_{t}\right)-\bm{\mu}_{t-1}\right),\\ \zeta_{t}\sim\mathcal{N}\left(\bm{\mu}_{t},\bm{C}_{t}\right),\\ \Theta_{t+1}\leftarrow\Theta_{t}-\eta\left(\hat{\bm{g}}_{s}\left(\Theta_{t}\right)+\psi\zeta_{t}\right).\end{gathered}$ (D.4)
In equation (D.4), we observe that the noise covariance pre-conditioner, scales the noise proportionally along the dimensions with larger gradients. This approach helps to escape saddle points and navigate the policy space more effectively while dealing with broader loss minima, ultimately leading to faster convergence and improved solutions. Therefore, incorporating adaptive noise into the loss function allows optimization of both the critic loss and the agent and adversarial policy loss functions. If $\alpha_{i}$ as the adversarial contribution, then the general action $a_{t}$ is expressed as follows:
$a_{t}=\alpha_{i}\nu_{\omega}\left(\Phi_{t}\right)+\left(1-\alpha_{i}\right)\pi_{\theta}\left(\Phi_{t}\right).$ (D.5)
@@
$y_{\textup{targ}}=r_{t}+\gamma\left(1-\mathfrak{d}\right)Q_{\phi_{\textup{targ}}}\left(\Phi^{\prime},\left(1-\alpha_{i}\right)\mu_{\theta}+\alpha_{i}\nu_{\omega}\right).$ (D.6)
@@
$L\left(\phi\right)=\frac{1}{N}\sum_{B=\left\{\left(\Phi_{t},a_{t},r_{t},{\Phi}^{\prime},\mathfrak{d}\right)\right\}}\left(y\left(r,\Phi^{\prime},d\right)-Q_{\phi}\left(\phi,a\right)\right)^{2}.$ (D.7)
@@
$\begin{gathered}\nabla_{\theta}\widehat{J\left(\theta,\omega_{t}\right)}=\frac{1-\alpha_{i}}{N}\sum_{\Phi\in\mathcal{D}}\nabla_{\theta}\mu_{\theta}\left(\Phi\right)\nabla_{a}Q_{\phi}\left(\Phi,a\right),\\ \nabla_{\omega}\widehat{J\left(\theta_{t},\omega\right)}=\frac{\alpha_{i}}{N}\sum_{\Phi\in\mathcal{D}}\nabla_{\omega}\nu_{\omega}\left(\Phi\right)\nabla_{a}Q_{\phi}\left(\Phi,a\right).\end{gathered}$ (D.8)
To minimize the loss functions of critics and policy given by equations (D.6) and (D.7) respectively, we incorporate SGLD noise into the update equations as specified in (D.4). This approach is taken to update the joint critic parameters $\phi$ and the agent and adversary policy parameters $\left\{\theta,\omega\right\}$. If we consider $\bm{g}=\left\{\nabla_{\theta}J,\nabla_{\omega}J\right\}$, $\bm{\mu}_{t}=\left\{\mu_{\theta},\mu_{\omega}\right\}$, $\bm{C}_{t}=\left\{C_{\theta},C_{\omega}\right\}$ according to the dynamics of SGLD given in (D.4), the update equation for the agent and the adversary parameter to obtain the mixed Nash equilibrium for the max-min objective in (19) for the $k^{\textup{th}}$ sample from the buffer at time $t$, is given as:
$\theta^{k+1}_{t}\leftarrow\theta^{k}_{t}+\eta\left(\nabla_{\theta}\widehat{J\left(\theta,\omega_{t}\right)}+\psi\xi_{t}\right),$ (D.9)
$\omega^{k+1}_{t}\leftarrow\omega^{k}_{t}-\eta\left(\nabla_{\omega}\widehat{J\left(\theta,\omega_{t}\right)}+\psi\xi_{t}\right).$ (D.10)
@@
$\bar{\omega}_{t}=\left(1-\beta\right)\bar{\omega}_{t}+\beta\omega^{\left(k+1\right)}_{t};\;\bar{\theta}_{t}=\left(1-\beta\right)\bar{\theta}_{t}+\beta\theta^{\left(k+1\right)}_{t}$. (D.11)
@@
$\omega_{t+1}=\left(1-\beta\right)\omega_{t}+\beta\bar{\omega}_{t};\;\theta_{t+1}=\left(1-\beta\right)\theta_{t}+\beta\bar{\theta}_{t}.$ (D.12)
@@
1:Observe state $\Phi$, select the actions according to $\rho_{\textup{agent}}=\mu_{\theta}+\zeta$, $\rho_{\textup{adv}}=\nu_{\omega}+{\zeta}^{\prime}$ where $\zeta$${\zeta}^{\prime}\sim\mathcal{N}\left(0,\sigma I\right)$.
@@
9:   Sample a random minibatch of NN transitions $B=\left\{\left(\Phi,\bar{\rho},r,{\Phi}^{\prime},d\right)\right\}$ from $\mathcal{D}$.
10:   Compute the targets as per (D.6).
11:   Compute the loss function to update the joint critic $\nabla_{\phi}L\left(\phi\right)$ as per (D.7).
12:   Compute the agent and adversary policy gradient according to (D.8).
13:   Compute the agent and adversary parameters as per (D.8).
14:   Update the parameters of the agent and adversary network $\left\{\theta,\omega\right\}$ according to (D.9) and (D.10).
15:   Obtain the value of $\left\{\bar{\theta},\bar{\omega}\right\}$ according to (D.11).
16:   Update the parameters of the target network used to compute the target according to (D.9).
$\begin{gathered}\phi_{\textup{targ}}\leftarrow\tau\phi_{\textup{targ}}+\left(1-\tau\right)\phi\\ \theta_{\textup{targ}}\leftarrow\tau\theta_{\textup{targ}}+\left(1-\tau\right)\theta^{k+1}_{t}\\ \omega_{\textup{targ}}\leftarrow\tau\omega_{\textup{targ}}+\left(1-\tau\right)\omega^{k+1}_{t}\end{gathered}$ (D.13)
@@
18:  Update $\left\{\omega_{t+1},\theta_{t+1}\right\}$ as per (D.12).
@@
22:Obtain $\Delta\mathbf{H}=\mathbb{E}\left[\Delta\mathbf{H}^{\textup{rand}}_{\alpha_{i}}-\max\left[\mathbf{H}^{\textup{opt}}_{\alpha_{i}}\right]\right]$.
@@
$\begin{gathered}\alpha_{i}\leftarrow\alpha_{i}+0.1,\\ \mathcal{P}^{\prime}\leftarrow\mathcal{P}\cup\left\{\theta,\omega\right\}_{\alpha_{i}},\\ \mathcal{Q}^{\prime}\leftarrow\mathcal{Q}\cup\left\{\phi\right\}_{\alpha_{i}},\\ \mathcal{P}\leftarrow\mathcal{P}^{\prime},\\ \mathcal{Q}^{\prime}\leftarrow\mathcal{Q}.\end{gathered}$ (D.14)
@@
Appendix E Algorithm for Generating Self-Induced Adversarial States
Algorithm 2 Generate self-induced adversarial state $\Phi^{\textup{adv}}$ for obtaining the distributional shift from vanilla and robust policies
@@ Appendix E Algorithm for Generating Self-Induced Adversarial States
4:  Compute the step size $\digamma=c\mathbin{/}k+c$.
@@
Appendix F Algorithm for Discounted Thompson Sampling
@@
Appendix G Antifragility Proof
$\lim_{t\to\infty}\mathbb{P}\left(\kappa_{t}=\arg\min_{k}d_{k}^{(t)}\right)=1.$ (G.1)
Let $k^{*}:=\arg\min_{k}d_{k}^{(t)}$ denote the optimal policy under the current (unseen) perturbation. Since DTS sampler utilizes the optimal policy $k^{*}$, and $r_{k}^{(t)}=d_{\min}^{(t)}/d_{k}^{(t)}$, we have:
$\lim_{t\to\infty}r_{\kappa_{t}}^{(t)}\to 1\quad\Rightarrow\quad\lim_{t\to\infty}d_{\kappa_{t}}^{(t)}\to d_{\min}^{(t)}.$ (G.2)
By the Lipschitz assumption:
$\mathbb{E}[R(\pi_{\kappa_{t}};\Phi_{t}^{\text{adv}})]\geq\mathbb{E}[R(\pi_{0};\Phi_{t})]-L\cdot d_{\kappa_{t}}^{(t)}.$ (G.3)
Differentiating both sides w.r.t. $t$, we get:
$\frac{d}{dt}\mathbb{E}[R(\pi_{\kappa_{t}};\Phi_{t}^{\text{adv}})]\geq-L\cdot\frac{d}{dt}d_{\kappa_{t}}^{(t)}.$ (G.4)
$\frac{d}{dt}d_{\kappa_{t}}^{(t)}<0\quad\Rightarrow\quad\frac{d}{dt}\mathbb{E}[R(\pi_{\kappa_{t}};\Phi_{t}^{\text{adv}})]>0.$ (G.5)
Appendix H Benchmarking Algorithms
Benchmarking is performed by evaluating the standard policy gradient algorithm, such as the DDPG, along with its robust variants. An approach involves an agent who intends to execute an action $\mathbf{a}$, but an alternative adversarial action $\bar{\mathbf{a}}$ is implemented with probability $\alpha$. Another variant, conceptually similar to the approach described in Section H.II, employs soft probabilistic robust policy iteration (PR-PI) as proposed in [tessler2019action], rather than solving the mixed Nash equilibrium objective in (19). The final benchmarking strategy involves training the agent to optimize the rewards without directly modifying the policy during training. Instead, an adversarial buffer is utilized, allowing the agent to learn the optimal value function in the presence of adversarial states. A brief explanation of these algorithms is provided in the subsequent subsections.
H.I Deep Deterministic Policy Gradient (DDPG)
The schematic shown in Figure 2, the action robust RL is replaced by DDPG considering $\alpha=0$. Hence, the critic target $y_{\textup{targ}}$ in (D.6) is modified as follows,
$y_{\textup{targ}}=r_{t}+\gamma\left(1-d\right)Q_{\phi_{\textup{targ}}}\left(\Phi^{\prime},\mu_{\theta}\right).$ (H.1)
Similarly the policy loss with $\alpha=0$ shown in (D.8) to compute the policy parameter $\theta$, will be as,
$\nabla_{\theta}\widehat{J\left(\theta\right)}=\frac{1}{N}\sum_{\Phi\in\mathcal{D}}\nabla_{\theta}\mu_{\theta}\left(\Phi\right)\nabla_{a}Q_{\phi}\left(\Phi,a\right).$ (H.2)
@@ H.I Deep Deterministic Policy Gradient (DDPG)
H.II Probabilistic Robust Markov Decision Process
@@
$\pi^{\textup{mix}}_{P,\alpha}\left(\left.\begin{matrix}\mathbf{a}\end{matrix}\right|\Phi\right)\equiv\left(1-\alpha\right)\pi\left(\left.\begin{matrix}\mathbf{a}\end{matrix}\right|\Phi\right)+\alpha\bar{\pi}\left(\left.\begin{matrix}\mathbf{a}\end{matrix}\right|\Phi\right).$ (H.3)
@@
H.III Noisy Robust Markov Decision Process
@@
$\pi^{\textup{mix}}_{N,\alpha}\left(\left.\begin{matrix}\mathbf{a}\end{matrix}\right|\Phi\right)=\mathbb{E}_{\mathbf{b}\sim\pi\left(.|\Phi\right),\bar{\mathbf{b}}\sim\bar{\pi}\left(.|\Phi\right)}\left[\vec{\mathbf{1}}_{\mathbf{a}=\left(1-\alpha\right)\mathbf{b}+\alpha\bar{\mathbf{b}}}\right].$ (H.4)
The NR-MDP is analogous to the Markov decision process (MDP) used to develop the ensembles, except for the addition of SGLD noise to the loss function, as described in equation (D.4) to obtain the mixed Nash equilibrium solution for the objective shown in (19). The solution to the min-max objective considered here is according to the soft PR-PI method as described in Algorithm 2 in [tessler2019action].
H.IV Learning with Adversarial Attacks